OpenStudy (mirandaregina):

If csc theta = √5, then what is cot theta?

Nnesha (nnesha):

what is the reciprocal of csc ?

OpenStudy (mirandaregina):

@Nnesha Sine

Nnesha (nnesha):

not just sin :=) 1/sin = csc so csc = ?

OpenStudy (anonymous):

In addition to sine, cosine, and tangent, there are three other trigonometric functions you need to know for the Math IIC: cosecant, secant, and cotangent. These functions are simply the reciprocals of sine, cosine, and tangent. Cotangent is the reciprocal of tangent.

OpenStudy (mirandaregina):

@Nnesha so csc = 1/sine

Nnesha (nnesha):

yes right so if csc = sqrt{5} then sin = ???

Nnesha (nnesha):

remember thy are reciprocal of each other

OpenStudy (mirandaregina):

@Nnesha then sin = 1/√5 = √5/5 ?

OpenStudy (phi):

draw a triangle and label its opposite side 1 and its hypotenuse sqr(5)

Nnesha (nnesha):

hmm \[\csc = \frac{ \sqrt{5} }{ 1 } ~~~~~ \sin=\frac{ 1 }{ \sqrt{5} }\] so draw right triangle for now it's okay leave the radical sign at the denominator  \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]

OpenStudy (mirandaregina):

@Nnesha so if sin theta = opp/hyp, and opp is 1 and hyp is √5, is it √5/5?

Nnesha (nnesha):

yes right sin = opp .hyp apply the Pythagorean theorem to find 3rd side of right triangle